A sample of 81 objects had the following results:

¯xx¯ = 197

The median = 199

SS = 10.

Suppose 5 was added to each of the numbers in the dataset.

What is the ¯xx¯ of the resulting dataset?

   

What is the median of the resulting dataset?

   

What is the SS of the resulting dataset?

   

What is the S2S2  of the resulting dataset
Answer #1

If 5 was added to each number in the dataset

xbar = 197 + 5 = 202

Median = 199 + 5 = 204

standard deviation, s = 10

variance, s^2 = 10*10 = 100